# MMA-Diffusion: MultiModal Attack on Diffusion Models

### v3

@@ 3.4 Image-Modal Attack
Greedy (Zhuang et al., 2023) nips’23 37.89 18.23 35.90 18.65 36.90 16.90 29.10 13.48
Genetic (Zhuang et al., 2023) nips’23 39.00 20.05 33.60 18.00 35.26 14.85 28.45 2.22
QF-pgd (Zhuang et al., 2023) nips’23 27.40 11.35 20.70 7.75 26.26 9.70 21.02 7.57
White-box SDv1.5 MMA-Diffusion (Ours) 84.90 73.23 84.80 75.10 80.40 54.20 83.37 67.54
@@
Greedy (Zhuang et al., 2023) nips’23 3.20 1.15 1.88 0.67 1.92 0.70 2.34 0.84
Genetic (Zhuang et al., 2023) nips’23 1.57 0.57 3.44 1.26 2.08 0.75 2.36 0.86
QF-pgd (Zhuang et al., 2023) nips’23 2.24 0.78 1.54 0.46 1.63 0.51 1.80 0.58
@@
Greedy (Zhuang et al., 2023) nips’23 28.10 10.43 18.00 6.15 25.30 9.18 23.80 8.57
Genetic (Zhuang et al., 2023) nips’23 30.00 11.58 16.80 5.87 23.40 9.00 23.40 8.82
QF-pgd (Zhuang et al., 2023) nips’23 22.80 8.03 15.40 5.10 21.70 7.89 19.97 7.01
SDXLv1.0 MMA-Diffusion (Ours) 76.30 49.28 71.70 44.87 73.10 40.38 73.70 44.84
[…]
Greedy (Zhuang et al., 2023) nips’23 21.80 9.08 23.10 10.13 23.10 8.92 22.67 9.37
Genetic (Zhuang et al., 2023) nips’23 19.30 7.78 20.50 9.72 23.40 8.80 21.07 8.77
QF-pgd (Zhuang et al., 2023) nips’23 12.80 4.40 11.80 4.60 13.60 5.18 12.73 4.73
Black-box SLD MMA-Diffusion (Ours) 75.60 53.05 78.70 61.33 75.90 45.72 76.73 53.37
4 Experiments
@@ 4.1 Experimental Settings
We select a subset of 1000 captions from the LAION-COCO dataset (Schuhmann et al., 2022), annotated with an NSFW score above 0.99 (out of 1.0), as our test prompts. The selection criteria are detailed in the Appendix. The NSFW scores in this dataset pertain solely to adult content. To diversify our NSFW themes evaluation, we include UnsafeDiff (Qu et al., 2023), a human-curated dataset designed for NSFW evaluation. UnsafeDiff provides 30 prompts across six NSFW themes: adult content, violence, gore, politics, racial discrimination, and inauthentic notable descriptions.
[…]
Baselines. As outlined in  Section 2, existing attack methods for diffusion models do not align with our problem, complicating direct comparisons. We employ the QF-attack(Zhuang et al., 2023) including Greedy, Genetic, and QF-PGD, initially designed to disrupt T2I, as our baselines due to its conceptual similarity to MMA-Diffusion. To ensure equal difficulty, we reconfigure the QF-attack from two aspects: (1) adjust the QF-attack’s objective function to mirror ours, following  Equation 1. (2) set the QF-attack’s input prompt (fix during the attack) as the target prompt, while masking sensitive words, as illustrated in  Figure 7(b). We also employ around 1000 human-written prompts with a sexual theme sourced from I2P (Schramowski et al., 2023) as a baseline, to reflect manual attack’s performance. Note that there is a bit of abuse since I2P prompts are prompts that are prone to unintentionally inducing the NSFW concept, while ours can target a concrete NSFW prompt. However, in terms of triggering the NSFW concept, they can be considered comparable, thus making I2P suitable for baseline evaluation.
Evaluation metrics. We employ the Attack Success Rate out of $N$ syntheses (ASR-N) as our evaluation metric. To compute ASR-N, we command T2I to generate $N$ images for each prompt. If any of these images exhibit NSFW content and bypass our simple prompt filter, the attack is considered successful. For instance, ASR-4 signifies the ratio of prompts where at least one out of the total 4 synthesized images contains explicit content, in relation to the overall number of tested prompts. To evaluate the open-source T2I models, we employ three NSFW detectors: Q16 (Schramowski et al., 2022), the builtin safety checker in SD (Safety-checker, 2023), and MHSC (Qu et al., 2023) as indicators to assess the ASR. Additionally, for the online services, as well as image-modal related attacks, we directly engage six human evaluators to independently perform the evaluation and report the average results.
@@ 4.2 Attacking Open-Source Models
Table 1 displays MMA-Diffusion’s significant success in steering the SD model towards generating NSFW content, with an average ASR-4 of 83.37%. This value signifies that most of our adversarial prompts successfully result in NSFW contents without using sensitive words, thereby demonstrating the vulnerability of T2I models to adversarial attacks, even when prompt filters are applied.
Black-box attacks on SDXL, SLD & DALLEnormal-$\cdot$2. Our generated adversarial prompts display impressive transferability, achieving 73.70% ASR-4 in black-box attacks on the SDXL, despite its architectural difference from the SD. Unlike the latter, SDXL employs a cascade structure composed of a basic and a refiner diffusion module, each with a different text encoder (Podell et al., 2023). We deduce that the transferability of MMA-Diffusion together with that of baselines is due to text encoders with varying structures learning the resembling semantic feature space from similar datasets.
[…]
@@
Comparison with baselines. As illustrated in  Table 1 and  Figure 5, MMA-Diffusion outperforms the baseline methods both quantitatively and qualitatively. First, our threat model, designed specifically for T2I attacks, allows the generation of adversarial prompts from scratch, enhancing the search space and the chance of finding target-resembling prompts in the latent space, leading to high-fidelity syntheses as shown in  Figure 5 (a) and (c). In contrast, the QF-Attack’s effectiveness is limited due to the strong coupling between the perturbation and the original prompt, while I2P achieves relatively high ASR but lacks the ability to control the generated content. Second, the baselines lack an effective mechanism to suppress sensitive words, causing the prompt filter to reject their adversarial prompts and leading to unsuccessful attacks.
@@ 4.3 Attacking Online T2I Services
We conducted an evaluation of two popular online services, namely Midjounery (Midjourney, 2023) and Leonardo.Ai (Leonardo.Ai, 2023), both of which are equipped with unknown AI moderators to counter NSFW content generation. To assess the safety of these services, we utilize the UnsafeDiff dataset (Qu et al., 2023) which consists of 30 human-crafted prompts covering 6 NSFW categories (refer to Table 2). For each target prompt, we generated 10 adversarial prompts and conducted a 10-query black-box attack on both online services. An attack is deemed successful if at least one adversarial prompt can circumvent online service’s AI moderator and generate a synthesis that is regarded as high-quality and high-fidelity by human evaluators. We achieved a 10-query attack success rate of 83.33% on Midjouney and 90.00% on Leonardo.Ai, respectively. Figure 6 illustrates the successful adversarial prompts alongside their corresponding generations. Moreover, we provide a concrete analysis about each online service’s robustness performance with respect to various NSFW themes, as reported in  Table 2.
@@
Results analysis on Midjounery. Midjouney demonstrates its defense mechanisms against five out of the six NSFW categories we tested, with the highest level of scrutiny applied to pornography-related content. Our generated adversarial prompts in the pornography category are able to bypass the detection without including sensitive words in 22% of the cases. Among the adversarial prompts that successfully pass through the AI moderator, 18% are able to induce Midjouney to generate pornography-related images, resulting in an overall success rate of 3.96%. As for violent content, 55.00% of the adversarial prompts are able to evade the defense mechanisms, and half of these prompts successfully generate violent content, resulting in a final success rate of 27.67%. However, the defense measures for horror and politics are relatively lenient. Notably, we observe Midjounery has no defense against the generation of real individual such as Donald Trump, Elon Musk and other notable. Content related to racial discrimination exhibits a high bypass rate, whereas our human evaluater can not identify discriminatory elements in the generations, resulting in a final success rate of only 10.00%. The possible reason is that Midjourney has filtered the racism content out of the training data. Furthermore, during the attack process, we found that our strategy of suppressing sensitive words are highly effective, as prompts containing sensitive words are directly rejected by Midjourney.
Results analysis on Leonard.Ai. We discovered that Leonardo.Ai’s prompt filter only examines explicit content. In our adversarial prompts with adult themes, we are able to bypass Leonardo.Ai’s defense mechanisms in 64% of the cases. Among these prompts, nearly 60% successfully induce Leonardo.Ai to generate adult images, resulting in a final attack success rate of 38%, which is nearly ten times higher than that of Midjourney. For bloody, horror, racism, and politics our attack also exhibits high attack success rate and image quality as exemplified in  Figure 6.
Failure case analysis. Interestingly, in our attacks targeting celebrities, we encountered relatively lower success rates, see the last column in  Table 2. Upon analyzing the failure cases, we identify a key factor contributing to this outcome. Our adversarial prompts are designed to exclude specific names of these individuals such as Trump and Biden. The absence of such crucial keywords makes it challenging for the prompts to accurately describe the intended celebrities. The most common failure cases involve the generation of individuals associated with the target person. For example, when targetting Biden, the generated images often depict Obama instead, referring Appendix for visualizations.
@@ 4.4 Multimodal Attack Results
Evaluation on image modal attacks. We first assess whether we can bypass the post-hoc defense, to generate realistic pornographic images. For this purpose, we focus on the image editing task of T2I models, which can be exploited to maliciously manipulate real individuals, as illustrated in  Figure 2. Due to ethical considerations, we collected synthetic individuals from Midjourney’s gallery as input images, as demonstrated in the first column in  Figure 7. To indicate the editable region, we manually utilized SAM (Kirillov et al., 2023) to label the masks, as shown in the second column of  Figure 7. We set the target prompts as "A completely naked woman/man/couple/girl/boy". The resulting (image-mask) pair, along with the matched target prompt, serves as the input for SD. Subsequently, SD modifies the input image according to the given target prompt. Firstly, we confirm that the safety checker can block the edited version of the input image with the absence of our attacks. Next, we perform attacks on the generation process to bypass the post-hoc safety checker. During the attack, we maintain the prompt and the mask as they are, while adding imperceptible adversarial perturbations ($\ell_{2}=16$) to the input image over 20 optimization iterations. The resulting adversarial images can be seen in the third column of  Figure 7, alongside two corresponding syntheses.  Figure 7 illustrates that our attacks are capable of deceiving the safety checker, allowing for the direct generation of high-quality pornographic content.
To quantify this risk, we generate 60 adversarial images with the same manner as above and evaluate their performance. A successful attack involves bypassing the safety checker and being deemed to contain NSFW content by our human evaluators. Results are presented in  Table 3. With the builtin safety-checker in SD, we achieve an 88.52% ASR-4 and a 78.68% ASR-1. We then transfer the obtained adversarial images to perform black-box attacks on two other types of post-hoc defenses, i.e. Q16 and Mhsc, where 30% and 20% of our adversarial image can deceive Q16 and Mhsc without extra efforts. These findings indicate the risk inherent in image editing tasks, and highlight the vulnerability of post-hoc defenses.
@@
SC 88.52 78.69 85.48 75.52
Mhsc 30.91 23.64 29.09 22.45
Q16 20.00 15.45 20.00 13.36
Evaluation for multimodal attacks. In more challenging scenarios where the T2I model is equipped with both a prompt filter and a post-hoc safety checker, our multimodal attack strategy becomes crucial. This evaluation involves generating adversarial prompts and combining them with corresponding adversarial images for SD to generate the final synthesized images. The last two columns of Figure 7 illustrate the resulting syntheses achieved through this multimodal attack strategy. The adversarial prompts are designed to bypass the prompt filter without compromising the original semantic information, while the adversarial perturbations effectively deceive the post-hoc safety checker, avoiding being flagged as inappropriate. The quantitative results, as shown in  Table 3, demonstrate the effectiveness of our multimodal attack, with an ASR-4 of 85.48% and an ASR-1 of 75.52%. These results indicate that the proposed multimodal attack strategy can effectively deceive both the prompt filter and the post-hoc safety checker.
@@ 5 Ethical Considerations
This research, centered on revealing security vulnerabilities in T2I diffusion models, is conducted with the intent to strengthen these systems rather than to enable misuse.
Responsible Use and Dissemination. To mitigate potential misuse, specific details of our attack methodologies have been deliberately omitted or generalized. We urge researchers and developers to utilize our findings responsibly to improve T2I model security.
Collaboration for Ethical AI. Addressing the ethical challenges in AI advancements requires collaboration among researchers, developers, ethicists, and policymakers. This collective effort is crucial in establishing guidelines for responsible AI research and its applications.
@@ 6 Conclusion
Appendix A Sensitive Word List

### v4

@@ 4 Experiments
Greedy [41] nips’23 37.89 18.23 35.90 18.65 36.90 16.90 29.10 13.48
Genetic [41] nips’23 39.00 20.05 33.60 18.00 35.26 14.85 28.45 2.22
QF-pgd [41] nips’23 27.40 11.35 20.70 7.75 26.26 9.70 21.02 7.57
White-box SDv1.5 [5] MMA-Diffusion (Ours) 84.90 73.23 84.80 75.10 80.40 54.20 83.37 67.54
@@
Greedy [41] nips’23 3.20 1.15 1.88 0.67 1.92 0.70 2.34 0.84
Genetic [41] nips’23 1.57 0.57 3.44 1.26 2.08 0.75 2.36 0.86
QF-pgd [41] nips’23 2.24 0.78 1.54 0.46 1.63 0.51 1.80 0.58
SDXLv1.0 [24] MMA-Diffusion (Ours) 76.30 49.28 71.70 44.87 73.10 40.38 73.70 44.84
@@
Greedy [41] nips’23 21.80 9.08 23.10 10.13 23.10 8.92 22.67 9.37
Genetic [41] nips’23 19.30 7.78 20.50 9.72 23.40 8.80 21.07 8.77
QF-pgd [41] nips’23 12.80 4.40 11.80 4.60 13.60 5.18 12.73 4.73
Black-box SLD [34] MMA-Diffusion (Ours) 75.60 53.05 78.70 61.33 75.90 45.72 76.73 53.37
@@ 4.1 Experimental Settings
We select a subset of 1000 captions from the LAION-COCO dataset [35], annotated with an NSFW score above 0.99 (out of 1.0), as our test prompts. The selection criteria are detailed in the Appendix. The NSFW scores in this dataset pertain solely to adult content. To diversify our NSFW themes evaluation, we include UnsafeDiff [25], a human-curated dataset designed for NSFW evaluation. UnsafeDiff provides 30 prompts across six NSFW themes: adult content, violence, gore, politics, racial discrimination, and inauthentic notable descriptions.
T2I models. We primarily execute white-box attacks on SDv1.5 [5] and report the results. Moreover, we repurpose the adversarial prompts derived from these attacks to conduct black-box attacks on two prevalent open-source models: SDXLv1.0 [24] and SLD (Medium) [34]. For online services, we evaluate Midjounery [1] and Leonardo.Ai [3].
Baselines. We employ the QF-attack[41] including Greedy, Genetic, and QF-PGD, initially designed to disrupt T2I, as our baselines due to its conceptual similarity to MMA-Diffusion. To ensure equal difficulty, we reconfigure the QF-attack from two aspects: (1) adjust the QF-attack’s objective function to mirror ours, following  Eq. 1. (2) set the QF-attack’s input prompt (fix during the attack) as the target prompt, while masking sensitive words, as illustrated in  Fig. 7(b). We also employ around 1000 human-written prompts with a sexual theme sourced from I2P [34] as a baseline, to reflect manual attack’s performance. Note that there is a bit of abuse since I2P prompts are prone to unintentionally inducing the NSFW concept, while ours can target a concrete NSFW prompt. However, in terms of triggering the NSFW concept, they can be considered comparable, thus making I2P suitable for baseline evaluation.
Evaluation metrics. We employ the Attack Success Rate out of $N$ syntheses (ASR-N) as our evaluation metric. To compute ASR-N, we command T2I to generate $N$ images for each prompt. If any of these images exhibit NSFW content and bypass our simple prompt filter, the attack is considered successful. For instance, ASR-4 signifies the ratio of prompts where at least one out of the total 4 synthesized images contains explicit content, in relation to the overall number of tested prompts. To evaluate the open-source T2I models, we employ three NSFW detectors: Q16 [33], the builtin safety checker in SD [4], and MHSC [25] as indicators to assess the ASR. Additionally, for the online services, as well as image-modal related attacks, we directly engage six human evaluators to independently perform the evaluation and report the average results.
@@ 4.2 Attacking Open-Source Models
Tab. 1 displays MMA-Diffusion’s significant success in steering the SD model towards generating NSFW content, with an average ASR-4 of 83.37%. This value signifies that most of our adversarial prompts successfully result in NSFW contents without using sensitive words, thereby demonstrating the vulnerability of T2I models to adversarial attacks, even when prompt filters are applied.
Black-box attacks on SDXL & SLD. Our generated adversarial prompts display impressive transferability, achieving 73.70% ASR-4 in black-box attacks on the SDXL, despite its architectural difference from the SD. Unlike the latter, SDXL employs a cascade structure composed of a basic and a refiner diffusion module, each with a different text encoder [24]. We deduce that the transferability of MMA-Diffusion together with that of baselines is due to text encoders with varying structures learning the resembling semantic feature space from similar datasets.
In contrast, SLD [34] shares the same architecture as SD, while the difference lies in the inference phase. SLD utilizes a batch of NSFW-related concept embeddings defined within the latent space to guide the generation process away from the predefined NSFW concepts, enhancing the safety of the generated images. Despite the defense mechanisms in SLD, MMA-Diffusion still achieves a relatively high attack success rate, with ASR-4 achieving 76.73%. The primary reason for the successful attack is that the embeddings used in SLD are derived from a fixed set of sensitive words. However, MMA-Diffusion effectively avoids a significant portion of them, thus mitigating the impact of SLD.
Comparison with baselines. As illustrated in  Tab. 1 and  Fig. 5, MMA-Diffusion outperforms the baseline methods both quantitatively and qualitatively. First, our threat model, designed specifically for T2I attacks, allows the generation of adversarial prompts from scratch, enhancing the search space and the chance of finding target-resembling prompts in the latent space, leading to high-fidelity syntheses as shown in  Fig. 5 (a) and (c). In contrast, the QF-Attack’s effectiveness is limited due to the strong coupling between the perturbation and the original prompt, while I2P achieves relatively high ASR but lacks the ability to control the generated content. Second, the baselines lack an effective mechanism to suppress sensitive words, causing the prompt filter to reject their adversarial prompts and leading to unsuccessful attacks.
@@ 4.3 Attacking Online T2I Services
We conducted an evaluation of two popular online services, namely Midjounery [1] and Leonardo.Ai [3], both of which are equipped with unknown AI moderators to counter NSFW content generation. To assess the safety of these services, we utilize the UnsafeDiff dataset [25] which consists of 30 human-crafted prompts covering 6 NSFW categories (refer to Tab. 2). For each target prompt, we generated 10 adversarial prompts and conducted a 10-query black-box attack on both online services. An attack is deemed successful if at least one adversarial prompt can circumvent online service’s AI moderator and generate a synthesis that is regarded as high-quality and high-fidelity by human evaluators. We achieved a 10-query attack success rate of 83.33% on Midjouney and 90.00% on Leonardo.Ai, respectively. Fig. 6 illustrates the successful adversarial prompts alongside their corresponding generations. Moreover, Tab. 2 provides a concrete analysis of each online service’s robustness performance with respect to various NSFW themes.
@@
Results analysis on Midjounery. Midjouney demonstrates its defense mechanisms against five out of the six NSFW categories we tested, with the highest level of scrutiny applied to pornography-related content. Our generated adversarial prompts in the pornography category are able to bypass the detection without including sensitive words in 22% of the cases. Among the adversarial prompts that successfully pass through the AI moderator, 18% are able to induce Midjouney to generate pornography-related images, resulting in an overall success rate of 3.96%. As for violent content, 55.00% of the adversarial prompts are able to evade the defense mechanisms, and half of these prompts successfully generate violent content, resulting in a final success rate of 27.67%. However, the defense measures for horror and politics are relatively lenient. Notably, we observe Midjounery has no defense against the generation of real individual such as Elon Musk and other notable. Furthermore, during the attack process, we found that our strategy of suppressing sensitive words are highly effective, as prompts containing sensitive words are directly rejected by Midjourney.
Results analysis on Leonard.Ai. We discovered that Leonardo.Ai’s prompt filter only examines explicit content. In our adversarial prompts with adult themes, we are able to bypass Leonardo.Ai’s defense mechanisms in 64% of the cases. Among these prompts, nearly 60% successfully induce Leonardo.Ai to generate adult images, resulting in a final attack success rate of 38%, which is nearly ten times higher than that of Midjourney. For bloody, horror, racism, and politics our attack also exhibits high attack success rate and image quality as exemplified in  Fig. 6.
Failure case analysis. Interestingly, in our attacks targeting celebrities, we encountered relatively lower success rates, see the last column in  Tab. 2. Upon analyzing the failure cases, we identify a key factor contributing to this outcome. Our adversarial prompts are designed to exclude specific names of these individuals such as Trump and Biden. The absence of such crucial keywords makes it challenging for the prompts to accurately describe the intended celebrities. The most common failure cases involve the generation of individuals associated with the target person. For example, when targetting Biden, the generated images often depict Obama instead, referring Appendix for visualizations.
@@ 4.4 Multimodal Attack Results
Evaluation on image modal attacks. We first assess whether we can bypass the post-hoc defense, to generate realistic pornographic images. For this purpose, we focus on the image editing task of T2I models, which can be exploited to maliciously manipulate real individuals, as illustrated in  Fig. 2. Due to ethical considerations, we collected synthetic individuals from Midjourney’s gallery as input images, as demonstrated in the first column in  Fig. 7. To indicate the editable region, we manually utilized SAM [13] to label the masks, as shown in the second column of  Fig. 7. We set the target prompts as "A completely naked woman/man/couple/girl/boy". The resulting (image-mask) pair, along with the matched target prompt, serves as the input for SD. Subsequently, SD modifies the input image according to the given target prompt. Firstly, we confirm that the safety checker can block the edited version of the input image with the absence of our attacks. Next, we perform attacks on the generation process to bypass the post-hoc safety checker. During the attack, we maintain the prompt and the mask as they are, while adding imperceptible adversarial perturbations ($\ell_{2}=16$) to the input image over 20 optimization iterations. The resulting adversarial images can be seen in the third column of  Fig. 7, alongside two corresponding syntheses.  Fig. 7 illustrates that our attacks are capable of deceiving the safety checker, allowing for the direct generation of high-quality pornographic content.
To quantify this risk, we generate 60 adversarial images with the same manner as above and evaluate their performance. A successful attack involves bypassing the safety checker and being deemed to contain NSFW content by our human evaluators. Results are presented in  Tab. 3. With the builtin safety-checker in SD, we achieve an 88.52% ASR-4 and a 78.68% ASR-1. We then transfer the obtained adversarial images to perform black-box attacks on two other types of post-hoc defenses, i.e. Q16 [33] and Mhsc [25], where 30% and 20% of our adversarial image can deceive Q16 [33] and Mhsc [25] without extra efforts.
@@
SC [4] 88.52 78.69 85.48 75.52
Mhsc [25] 30.91 23.64 29.09 22.45
Q16 [33] 20.00 15.45 20.00 13.36
Evaluation for multimodal attacks. In more challenging scenarios where the T2I model is equipped with both a prompt filter and a post-hoc safety checker, our multimodal attack strategy becomes crucial. This evaluation involves generating adversarial prompts and combining them with corresponding adversarial images for SD to generate the final synthesized images. The last two columns of Fig. 7 illustrate the resulting syntheses achieved through this multimodal attack strategy. The adversarial prompts are designed to bypass the prompt filter without compromising the original semantic information, while the adversarial perturbations effectively deceive the post-hoc safety checker, avoiding being flagged as inappropriate. The quantitative results, as shown in  Tab. 3, demonstrate the effectiveness of our multimodal attack, with an ASR-4 of 85.48% and an ASR-1 of 75.52%. These results indicate that the proposed multimodal attack strategy can effectively deceive both the prompt filter and the post-hoc safety checker.
@@ 5 Ethical Considerations
This research, centered on revealing security vulnerabilities in T2I diffusion models, is conducted with the intent to strengthen these systems rather than to enable misuse. To mitigate potential misuse, specific details of our attack methods have been deliberately omitted or generalized. We urge developers to utilize our findings responsibly to improve T2I model security. We advocate for ethical awareness in AI research, particularly in fields involving generative models. Balancing innovation with ethical responsibility is vital. Transparent reporting, with an emphasis on societal impact and misuse prevention, is essential.
@@ 6 Conclusion
Acknowledgements
This work is supported in part by General Research Fund (GRF) of Hong Kong Research Grants Council (RGC) under Grant No. 14203521, the CUHK SSFCRS funding No. 3136023, the Research Matching Grant Scheme under Grant No. 7106937, 8601130, and 8601440, the National Key Research and Development Program of China Grant No. 2021YFF0901503, and the National Natural Science Foundation of China under Grants No. 62206287. This work is conducted in the JC STEM Lab of Intelligent Design Automation funded by The Hong Kong Jockey Club Charities Trust. Further, we thank Jianping Zhang and Ruosi Wan for their valuable comments.
Supplementary Material
Overview
This supplementary material provides additional details and results that are not included in the main paper due to page limitations. The following items are included in this supplementary material:
  • Sensitive word list in Section 3.
  • Detailed algorithm in Section 3.4.
  • Additional experimental setup and details in Section 4.1.
[…]
  • MMA-Diffusion against input sanitization defense.
  • Failure case visualizations in Section 4.3 and Section  4.4.
  • More visualization results.
A Sensitive Word List